# Rethinking Impersonation and Dodging Attacks on Face Recognition Systems

### v3

@@ Abstract
Face Recognition (FR) systems can be easily deceived by adversarial examples that manipulate benign face images through imperceptible perturbations. Adversarial attacks on FR encompass two types: impersonation (targeted) attacks and dodging (untargeted) attacks. Previous methods often achieve a successful impersonation attack on FR; However, it does not necessarily guarantee a successful dodging attack on FR in the black-box setting. In this paper, our key insight is that the generation of adversarial examples should perform both impersonation and dodging attacks simultaneously. To this end, we propose a novel attack method termed as Adversarial Pruning (Adv-Pruning), to fine-tune existing adversarial examples to enhance their dodging capabilities while preserving their impersonation capabilities. Adv-Pruning consists of Priming, Pruning, and Restoration stages. Concretely, we propose Adversarial Priority Quantification to measure the region-wise priority of original adversarial perturbations, identifying and releasing those with minimal impact on absolute model output variances. Then, Biased Gradient Adaptation is presented to adapt the adversarial examples to traverse the decision boundaries of both the attacker and victim by adding perturbations favoring dodging attacks on the vacated regions, preserving the prioritized features of the original perturbations while boosting dodging performance. As a result, we can maintain the impersonation capabilities of original adversarial examples while effectively enhancing dodging capabilities. Comprehensive experiments demonstrate the superiority of our method compared with state-of-the-art adversarial attacks.
1 Introduction
@@
Figure 1: Top: previous methods that achieve a successful impersonation attack on FR cannot guarantee a successful dodging attack on FR in the black-box setting. In contrast, we present Adv-Pruning, including Priming, Pruning, and Restoration Stages, to perform both impersonation and dodging attacks simultaneously. Bottom (left): natural Multi-identity Samples (MS). Bottom (right): the dodging Attack Success Rate (%) between the previous methods and Adv-Pruning on multiple models.
@@
Thanks to the ceaseless advancements in deep learning, Face Recognition (FR) has achieved exceptional performance [41; 49; 9; 1; 2; 26]. However, the vulnerability of existing FR models to adversarial attacks poses a significant threat to their security. Hence, there is an urgent need to enhance the performance of adversarial face examples to expose more blind spots in FR models. As a result, several research endeavors have been directed towards this realm. A multitude of adversarial attacks have been developed to create adversarial face examples with characteristics such as stealthiness [40; 62; 7; 19; 42], transferability [70; 29; 72; 71], and physical attack capability [64; 63; 28]. These efforts contribute to enhancing the effectiveness of adversarial attacks on FR. Nevertheless, these studies primarily concentrate on bolstering either impersonation attacks or dodging attacks, overlooking the exploration of the effectiveness of dodging attacks when crafting adversarial face examples using impersonation attacks.
@@
In real-world deployment contexts, individuals with malicious intent are prone to creating adversarial face examples incorporating their own facial features to manipulate FR systems to mistakenly identify them as pre-defined victims during impersonation attacks. Concurrently, the individuals strive to evade accurate identification as perpetrators, thereby circumventing detection and preventing legal accountability. This requires the creation of adversarial examples capable of executing both impersonation and dodging attacks simultaneously. In the realm of adversarial attacks on image classification, a successful impersonation attack typically implies a successful dodging attack. However, FR is an open-set task [49; 9], which is quite different from image classification. In the real-world deployment of FR systems, accurately predicting the class probability of identities presents an extreme challenge. Therefore, we extract embeddings from two face images using the FR model. Subsequently, the distance between the two embeddings is used to determine whether the images belong to the same identity. If the distance falls below a predefined threshold, the two images are recognized as belonging to the same identity; otherwise, they are classified as different identities. Based on the measurement of FR, there are two decision boundaries for each FR model when crafting adversarial examples as shown in Fig. 1. As a result, there exists the natural samples that can be classified as two different identities in theory. We denote these samples as multi-identity samples (refer to Section 4.4).
@@
The existence of multi-identity samples implies that a successful impersonation attack on FR does not necessarily guarantee a successful dodging attack on FR. Existing research indicates that an adversarial sample was located near the decision boundary [18; 4]. Suppose we generate adversarial face examples using previous methods. In the white-box setting, both the structures and parameters of the victim models are known, enabling the generation of adversarial face examples that can cross the decision boundaries of both attacker and victim, as shown in Fig. 1. However, in the black-box setting, the decision boundaries of black-box models differ from those of the surrogate models. Consequently, adversarial examples generated on the surrogate model lie near the decision boundary of the victim, preventing them from crossing the decision boundary of the attacker. As such, the majority of adversarial face examples crafted by previous methods, which can successfully perform impersonation attacks, fail to perform dodging attacks in the black-box setting.
Figure 2: Overview of our Adv-Pruning attack framework, which consists of Priming, Pruning, and Restoration stages. (a) During the Priming stage, we optimize the adversarial examples to ensure they have sufficient attack performance. (b) In the Pruning stage, we propose Adversarial Priority Quantification to quantify the priority of adversarial perturbations. Subsequently, we sparsify the adversarial perturbations based on the quantified priorities. (c) In the Restoration stage, we present Biased Gradient Adaptation to introduces gradient perturbations biased to dodging attacks on the sparsified regions.
@@
In this paper, we propose a novel attack method, termed as Adversarial Pruning (Adv-Pruning). In the realm of adversarial attacks on FR, previous impersonation methods have achieved a significant level of sophistication. However, there remains a pressing need to bolster the efficacy of adversarial face examples in dodging attacks. Consequently, our research is directed towards enhancing the dodging attack performance of adversarial face examples while maintaining the impersonation attack performance. Specifically, we introduce an attack consisting of three stages: Priming, Pruning, and Restoration. In the Priming stage, we optimize the adversarial examples to ensure adequate attack potential. In the Pruning stage, with considering the pruning concept in model compression, we propose Adversarial Priority Quantification to measure the region-wise priority of original adversarial perturbations using an priority measure which is directly proportional to the supremum of the absolute model output variances. After processing by Adversarial Priority Quantification, we prune the adversarial face examples to free up less prioritized adversarial perturbations. In the Restoration stage, we propose Biased Gradient Adaptation to add biased gradient perturbations favoring dodging attacks on the pruned regions to adapt the adversarial face examples into the space that can be classified as the victim while remaining unidentifiable as the attacker, thereby enhancing the dodging performance of the adversarial face examples without compromising the prioritized features of original adversarial perturbations. As illustrated in the top of Fig. 1, after undergoing these stages, the adversarial face example generated by our proposed method can successfully traverse the decision boundaries of both the attacker and victim of the black-box model, achieving successful black-box impersonation and dodging attacks.
@@
2 Related Work
2.1 Adversarial Attacks
@@
The primary objective of adversarial attacks is to introduce imperceptible perturbations to benign images to deceive machine learning systems and cause them to make mistakes [47; 14]. The existence of adversarial examples poses a significant threat to the security of current machine learning systems. Lots of efforts have been dedicated to researching adversarial attacks in order to enhance the robustness of these systems [33; 69; 30; 55; 34; 43; 12; 73; 68; 37]. To improve the performance of black-box adversarial attacks, DI [60] applies random transformations to adversarial examples in each iteration to achieve a data augmentation effect. VMI-FGSM [53] employs gradient variance to stabilize the updating process of adversarial examples, boosting the black-box performance. SSA [33] transforms adversarial examples into the frequency domain and uses spectrum transformation to augment them. SIA [54] applies a random image transformation to each image block, generating a varied collection of images that are then employed for gradient calculation. BSR [51] divides the input image into multiple blocks, subsequently shuffling and rotating these blocks in a random manner, creating a collection of new images for the purpose of gradient calculation. DA [16] utilizes dispersion amplification to enhance the multi-task attack capability of adversarial attacks. Despite their gratifying progress, these studies neglect the consideration of pruning adversarial examples through introducing pruning methods into the realm of adversarial attacks. In our research, we propose a novel pruning method capable of identifying and freeing up the adversarial perturbations with minimal impact on absolute model output variances, thereby sparsifying regions for adding adversarial perturbations with the aim of dodging capabilities improvement.
2.2 Adversarial Attacks on Face Recognition
@@
Based on the restriction of the adversarial perturbations, adversarial attacks on FR can be classified into two categories: restricted attacks [32; 61; 10; 74; 35; 31; 5] and unrestricted attacks [3; 6; 65; 8; 44; 46; 48; 56; 57]. Restricted attacks on FR are the attacks that generate adversarial examples in a restricted bound (e.g. $L_{p}$ bound). To enhance the transferability of adversarial attacks on FR, [70] propose DFANet, which applies dropout on the feature maps of the convolutional layers to achieve ensemble-like effects. In addition, [72] introduces BPFA, which further improves the transferability of adversarial attacks on FR by incorporating beneficial perturbations [58] on the feature maps of the FR models, resulting in hard model augmentation effects. [29] leverages extra information from FR-correlated tasks and uses a multi-task optimization framework to enhance the transferability of crafted adversarial examples. The unrestricted adversarial attacks on FR are the attacks that generate adversarial examples without the restriction of a predefined perturbation bound. They mainly focus on physical attacks [59; 63; 28], attribute editing [40; 23] and generating adversarial examples based on makeup transfer [64; 19; 42]. The existing literature on both restricted and unrestricted adversarial attacks on FR systems has successfully enhanced the performance of these attacks. Nevertheless, it remains under-explored in the correlation between impersonation and dodging attacks. This paper elegantly addresses this by investigating the correlation between impersonation and dodging attacks and introducing a novel attack method that bolsters the dodging capabilities while preserving the impersonation capabilities of previous methods.
3 Methodology
3.1 Problem Formulation
@@
Let $\mathcal{F}^{vct}(x)$ denote the FR model used by the victim to extract the embedding from a face image $x$. We refer to $x^{s}$ and $x^{t}$ as the attacker and victim images, respectively. The objective of the impersonation attacks explored in our research is to manipulate $\mathcal{F}^{vct}$ in order to misclassify $x^{adv}$ as $x^{t}$, while ensuring that $x^{adv}$ bears a close visual resemblance to $x^{s}$. By contrast, the objective of the dodging attacks proposed in this study is to render $\mathcal{F}^{vct}(x)$ unable to identify $x^{adv}$ as $x^{s}$, while simultaneously ensuring that $x^{adv}$ bears a visual resemblance to $x^{s}$. For the sake of clarity and conciseness, the detailed optimization objectives for both impersonation and dodging attacks are provided in the appendices.
@@
Few works explore the correlation between impersonation and dodging attacks on FR. In the following, we delve into the correlation between these two types of attacks and propose a novel method to enhance dodging attacks while maintaining impersonation attacks. An overview of the proposed method is illustrated in Fig. 2. As depicted in Fig. 2, our proposed method is structured into three stages: Priming, Pruning, and Restoration. Through the sequential application of these stages, we are able to generate adversarial examples that exhibit a potent combination of impersonation and dodging attack capabilities.
3.2 Exploring the Impersonation and Dodging Attack on Face Recognition
@@
In most cases, the victim model $\mathcal{F}^{vct}$ is not accessible to the attacker, making it extremely challenging to optimize the objectives for black-box attacks directly. To circumvent this issue, a common approach is to leverage a surrogate model $\mathcal{F}$ accessible to the attacker to generate adversarial examples that can be transferred to the victim model for an effective attack [66; 69; 39; 54; 51; 38; 11; 13; 27; 5].
@@
$\mathcal{L}^{i}=\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{t}\right)\right)\|^{2}_{2}$ (1)
@@
$\mathcal{L}^{d}=-\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{s}\right)\right)\|^{2}_{2}$ (2)
As the FR task is an open-set task, it is impractical to predict the classes of users during the practical deployment of the FR model. Therefore, we need to compare the distance between two face images to discern whether they depict the same identity or not. Based on the identification method in FR, multi-identity samples exist theoretically. Our experiments verify the existence of such samples among benign face images. The existence of multi-identity samples raises a question:
@@
To this end, we generate adversarial face examples using the previous impersonation attack and evaluate its dodging Attack Success Rate (ASR). Our experiment confirms that the majority of adversarial examples crafted through previous methods, which are successful in performing impersonation attacks, fail to successfully execute dodging attacks in the black-box setting (see Section 4.4).
@@
To accomplish this objective, a straightforward approach is to generate adversarial face examples using a multi-task attack strategy. In the following, we will take the Lagrangian attack strategy as the example for its simplicity. The Lagrangian attack strategy utilizes the following loss function to craft adversarial examples:
$\mathcal{L}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$ (3)
@@
3.3 Adversarial Pruning Attack

### v4

@@ Abstract.
Face Recognition (FR) systems can be easily deceived by adversarial examples that manipulate benign face images through imperceptible perturbations. Adversarial attacks on FR encompass two types: impersonation (targeted) attacks and dodging (untargeted) attacks. Previous methods often achieve a successful impersonation attack on FR, however, it does not necessarily guarantee a successful dodging attack on FR in the black-box setting. In this paper, our key insight is that the generation of adversarial examples should perform both impersonation and dodging attacks simultaneously. To this end, we propose a novel attack method termed as Adversarial Pruning (Adv-Pruning), to fine-tune existing adversarial examples to enhance their dodging capabilities while preserving their impersonation capabilities. Adv-Pruning consists of Priming, Pruning, and Restoration stages. Concretely, we propose Adversarial Priority Quantification to measure the region-wise priority of original adversarial perturbations, identifying and releasing those with minimal impact on absolute model output variances. Then, Biased Gradient Adaptation is presented to adapt the adversarial examples to traverse the decision boundaries of both the attacker and victim by adding perturbations favoring dodging attacks on the vacated regions, preserving the prioritized features of the original perturbations while boosting dodging performance. As a result, we can maintain the impersonation capabilities of original adversarial examples while effectively enhancing dodging capabilities. Comprehensive experiments demonstrate the superiority of our method compared with state-of-the-art adversarial attack methods.
Face Recognition, Adversarial Attacks, Adversarial Attacks on Face Recognition, Impersonation Attacks, Dodging Attacks
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3664647.3681440conference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10submissionid: 430ccs: Computing methodologies Biometrics
1. Introduction
@@
Figure 1. Top: previous methods that achieve a successful impersonation attack on FR cannot guarantee a successful dodging attack on FR in the black-box setting. In contrast, we present Adv-Pruning, including Priming, Pruning, and Restoration Stages, to perform both impersonation and dodging attacks simultaneously. Bottom (left): benign Multi-identity Samples (MS). Bottom (right): the dodging Attack Success Rate (%) between the previous methods and Adv-Pruning on multiple models.
@@
Thanks to the ceaseless advancements in deep learning, Face Recognition (FR) has achieved exceptional performance (Schroff et al., 2015; Wang et al., 2018; Deng et al., 2022; An et al., 2021; Boutros et al., 2023; Li et al., 2023a). However, the vulnerability of existing FR models poses a significant threat to their security (Das et al., 2021; Zhou et al., 2022a, b, 2023c; Narayan et al., 2023; Wang et al., 2024b; Zhou et al., 2024b), with adversarial attacks being one of the key concerns. Hence, there is an urgent need to enhance the performance of adversarial face examples to expose more blind spots in FR models. As a result, several research endeavors have been directed towards this realm. A multitude of adversarial attacks have been developed to create adversarial face examples with characteristics such as stealthiness (Qiu et al., 2020; Yang et al., 2021; Cherepanova et al., 2021; Hu et al., 2022; Shamshad et al., 2023), transferability (Zhong and Deng, 2021; Li et al., 2023d; Zhou et al., 2023b, 2024a), and physical attack capability (Yin et al., 2021; Yang et al., 2023; Li et al., 2023c). These efforts contribute to enhancing the effectiveness of adversarial attacks on FR. Nevertheless, these studies primarily concentrate on bolstering either impersonation attacks or dodging attacks, overlooking the exploration of the effectiveness of dodging attacks when crafting adversarial face examples using impersonation attacks.
@@
In real-world deployment contexts, individuals with malicious intent are prone to creating adversarial face examples incorporating their own facial features to manipulate FR systems to mistakenly identify them as pre-defined victims during impersonation attacks. Concurrently, the individuals strive to evade accurate identification as perpetrators, thereby circumventing detection and preventing legal accountability. This requires the creation of adversarial examples capable of executing both impersonation and dodging attacks simultaneously. In the realm of adversarial attacks on image classification, a successful impersonation attack typically implies a successful dodging attack. However, FR is an open-set task (Wang et al., 2018; Deng et al., 2022), which is quite different from image classification. In the real-world deployment of FR systems, accurately predicting the class probability of identities presents an extreme challenge. Therefore, we extract embeddings from two face images using the FR model. Subsequently, the distance between the two embeddings is used to determine whether the images belong to the same identity. If the distance falls below a predefined threshold, the two images are recognized as belonging to the same identity; otherwise, they are classified as different identities. Based on the measurement of FR, there are two decision boundaries for each FR model when crafting adversarial examples as shown in Figure 1. As a result, there exists the benign samples that can be classified as two different identities in theory. We denote these samples as multi-identity samples.
@@
The existence of multi-identity samples implies that a successful impersonation attack on FR does not necessarily guarantee a successful dodging attack on FR. Existing research indicates that an adversarial sample was located near the decision boundary (Heo et al., 2019; Cao and Gong, 2017). Suppose we generate adversarial face examples using previous methods. In the white-box setting, both the structures and parameters of the victim models are known, enabling the generation of adversarial face examples that can cross the decision boundaries of both attacker and victim, as shown in Figure 1. However, in the black-box setting, the decision boundaries of black-box models differ from those of the surrogate models. Consequently, adversarial examples generated on the surrogate model lie near the decision boundary of the victim, preventing them from crossing the decision boundary of the attacker. As such, the majority of adversarial face examples crafted by previous methods, which can successfully perform impersonation attacks, fail to perform dodging attacks in the black-box setting.
@@
In this paper, we propose a novel attack method, termed as Adversarial Pruning (Adv-Pruning). In the realm of adversarial attacks on FR, previous impersonation methods have achieved a significant level of sophistication. However, there remains a pressing need to bolster the efficacy of adversarial face examples in dodging attacks. Consequently, our research is directed towards enhancing the dodging attack performance of adversarial face examples while maintaining the impersonation attack performance. Specifically, we introduce an attack consisting of three stages: Priming, Pruning, and Restoration. In the Priming stage, we optimize the adversarial examples to ensure adequate attack potential. In the Pruning stage, considering the pruning concept in model compression, we propose Adversarial Priority Quantification to measure the region-wise priority of original adversarial perturbations using a priority measure that is directly proportional to the supremum of the absolute model output variances. After processing by Adversarial Priority Quantification, we prune the adversarial face examples to free up less prioritized adversarial perturbations. In the Restoration stage, we propose Biased Gradient Adaptation to add biased gradient perturbations favoring dodging attacks on the pruned regions to adapt the adversarial face examples into the space that can be classified as the victim while remaining unidentifiable as the attacker, thereby enhancing the dodging performance of the adversarial face examples without compromising the prioritized features of original adversarial perturbations. As illustrated in the top of Figure 1, after undergoing these stages, adversarial face examples generated by our method can successfully traverse the decision boundaries of both the attacker and victim of the black-box model, achieving successful black-box impersonation and dodging attacks.
@@
2. Related Work
2.1. Adversarial Attacks
@@
The primary objective of adversarial attacks is to introduce imperceptible perturbations to benign images to deceive machine learning systems and cause them to make mistakes (Szegedy et al., 2014; Goodfellow et al., 2015). The existence of adversarial examples poses a significant threat to the security of current machine learning systems. Lots of efforts have been dedicated to researching adversarial attacks in order to enhance the robustness of these systems (Long et al., 2022; Liang et al., 2023; Lu et al., 2023; Shayegani et al., 2024; Ge et al., 2023a; Zhou et al., 2023a; Zhang et al., 2022b; Mingxing et al., 2021). To improve the performance of black-box adversarial attacks, DI (Xie et al., 2019) applies random transformations to adversarial examples in each iteration to achieve a data augmentation effect. VMI-FGSM (Wang and He, 2021) employs gradient variance to stabilize the updating process of adversarial examples, boosting the black-box performance. SSA (Long et al., 2022) transforms adversarial examples into the frequency domain and uses spectrum transformation to augment them. SIA (Wang et al., 2023) applies a random image transformation to each image block, generating a varied collection of images that are then employed for gradient calculation. BSR (Wang et al., 2024a) divides the input image into multiple blocks, subsequently shuffling and rotating these blocks in a random manner, creating a collection of new images for the purpose of gradient calculation. DA (Haleta et al., 2021) utilizes dispersion amplification to enhance the multi-task attack capability of adversarial attacks. Despite their gratifying progress, these studies neglect the consideration of pruning adversarial examples by introducing pruning methods into the realm of adversarial attacks. In our research, we propose a novel pruning method capable of identifying and freeing up the adversarial perturbations with minimal impact on absolute model output variances, thereby sparsifying regions for adding adversarial perturbations with the aim of dodging capabilities improvement.
2.2. Adversarial Attacks on Face Recognition
@@
Based on the restriction of the adversarial perturbations, adversarial attacks on FR can be classified into two categories: restricted attacks and unrestricted attacks. Restricted attacks on FR are the attacks that generate adversarial examples in a restricted bound (e.g. $L_{p}$ bound). To enhance the transferability of adversarial attacks on FR, Zhong and Deng (Zhong and Deng, 2021) propose DFANet, which applies dropout on the feature maps of the convolutional layers to achieve ensemble-like effects. In addition, Zhou et al. (Zhou et al., 2023b) introduces BPFA, which further improves the transferability of adversarial attacks on FR by incorporating beneficial perturbations (Wen and Itti, 2020) on the feature maps of the FR models, resulting in hard model augmentation effects. Li et al. (Li et al., 2023d) leverages extra information from FR-related tasks and uses a multi-task optimization framework to enhance the transferability of crafted adversarial examples. Zhou el al. (Zhou et al., 2024c) propose an adversarial attack that can attack both FR and Face Anti-spoofing (FAS) models simultaneously, aiming to enhance the practicability of adversarial attacks on FR systems. The unrestricted adversarial attacks on FR are the attacks that generate adversarial examples without the restriction of a predefined perturbation bound. They mainly focus on physical attacks (Xiao et al., 2021; Yang et al., 2023; Li et al., 2023c), attribute editing (Qiu et al., 2020; Jia et al., 2022) and generating adversarial examples based on makeup transfer (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023). The existing literature on both restricted and unrestricted adversarial attacks on FR systems has successfully enhanced the performance of these attacks. Nevertheless, it remains under-explored in the correlation between impersonation and dodging attacks on FR. This paper addresses this by investigating the correlation between impersonation and dodging attacks and introducing a novel attack method that bolsters the dodging capabilities while preserving the impersonation capabilities of previous methods.
3. Methodology
3.1. Problem Formulation
@@
Let $\mathcal{F}^{vct}(x)$ denote the FR model used by the victim to extract the embedding from a face image $x$. We refer to $x^{s}$ and $x^{t}$ as the attacker and victim images, respectively. The objective of the impersonation attacks explored in our research is to manipulate $\mathcal{F}^{vct}$ in order to misclassify $x^{adv}$ as $x^{t}$, while ensuring that $x^{adv}$ bears a close visual resemblance to $x^{s}$. By contrast, the objective of the dodging attacks proposed in this study is to render $\mathcal{F}^{vct}(x)$ unable to identify $x^{adv}$ as $x^{s}$, while simultaneously ensuring that $x^{adv}$ bears a visual resemblance to $x^{s}$. For the sake of clarity and conciseness, the detailed optimization objectives for both impersonation and dodging attacks are provided in the supplementary.
@@
Few works explore the correlation between impersonation and dodging attacks on FR. In the following, we delve into the correlation between these two types of attacks and propose a novel method to enhance dodging attacks while maintaining impersonation attacks. An overview of the proposed method is illustrated in Figure 2. As depicted in Figure 2, our proposed method is structured into three stages: Priming, Pruning, and Restoration. Through the sequential application of these stages, we are able to generate adversarial examples that exhibit a potent combination of impersonation and dodging attack capabilities.
Figure 2. Overview of our Adv-Pruning attack framework, which consists of Priming, Pruning, and Restoration stages. (a) During the Priming stage, we optimize the adversarial examples to ensure they have sufficient attack performance. (b) In the Pruning stage, we propose Adversarial Priority Quantification to quantify the priority of adversarial perturbations. Subsequently, we sparsify the adversarial perturbations based on the quantified priorities. (c) In the Restoration stage, we present Biased Gradient Adaptation to introduce gradient perturbations biased to dodging attacks on the sparsified regions.
3.2. Exploring the Impersonation and Dodging Attack on Face Recognition
@@
In most cases, the victim model $\mathcal{F}^{vct}$ is not accessible to the attacker, making it extremely challenging to optimize the objectives for black-box attacks directly. To circumvent this issue, a common approach is to leverage a surrogate model $\mathcal{F}$ accessible to the attacker to generate adversarial examples that can be transferred to the victim model for an effective attack (Yuan et al., 2022; Zhang et al., 2022a; Naseer et al., 2023; Wang et al., 2023, 2024a; kanth Nakka and Salzmann, 2021; Gao et al., 2023; Ge et al., 2023b; Li et al., 2023b; Chen et al., 2023).
@@
(1) $\mathcal{L}^{i}=\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{t}\right)\right)\|^{2}_{2}$
@@
(2) $\mathcal{L}^{d}=-\|\phi\left(\mathcal{F}\left(x^{adv}\right)\right)-\phi\left(\mathcal{F}\left(x^{s}\right)\right)\|^{2}_{2}$
As the FR task is an open-set task, it is impractical to predict the classes of users during the practical deployment of the FR model. Therefore, we need to compare the distance between two face images to discern whether they depict the same identity or not. Based on the identification method in FR, multi-identity samples exist theoretically. Our experiments verify the existence of such samples among benign face images (see the supplementary). The existence of multi-identity samples raises a question:
@@
To this end, we generate adversarial face examples using the previous impersonation attack and evaluate its dodging Attack Success Rate (ASR). Our experiment confirms that the majority of adversarial examples crafted through previous methods, which are successful in performing impersonation attacks, fail to successfully execute dodging attacks in the black-box setting.
@@
To accomplish this objective, a straightforward approach is to generate adversarial face examples using a multi-task attack strategy. In the following, we will take the Lagrangian attack strategy as an example for its simplicity. The Lagrangian attack strategy utilizes the following loss function to craft adversarial examples:
(3) $\mathcal{L}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$
@@
3.3. Adversarial Pruning Attack